OpenStudy (anonymous):

Prove or disprove that the given point lies on the given circle. Point (4,-4) with center (1,0) and containing the point (1,5) THANK YOU :)

OpenStudy (phi):

circle with center (1,0) and containing the point (1,5) I would use the equation of a circle: \[ (x-h)^2 + (y-k)^2 = r^2 \] where the center is (h,k) and the radius is r we know the center is (1,0) so we can write \[ (x-1)^2 + y^2 = r^2 \] (notice we can write (y-0)^2 as just y^2 ) to find the radius, use the distance formula to find the distance from the center (1,0) to the point (1,5) we will find the distance is 5, so r=5, and the full equation is \[ (x-1)^2 + y^2 = 25 \] (remember that r^2 means r*r , and r is 5, so 5*5 or 25) now test the point (4, -4) in that equation: \[ (4-1)^2 + (-4)^2 =? 25 \] will we get 25 ? if we do, then (4, -4) is on the circle.